Flexible Budgeting and Variance Analysis

I Love My Chocolate Company makes dark chocolate and light chocolate. Both products require cocoa and sugar. The following planning information has been made available:

Standard Amount per Case
     Dark Chocolate      Light Chocolate      Standard Price per Pound
Cocoa 12 lbs. 9 lbs. $4.30
Sugar 10 lbs. 14 lbs. 0.60
Standard labor time 0.4 hr. 0.5 hr.
Dark Chocolate Light Chocolate
Planned production 4,200 cases 13,800 cases
Standard labor rate $16.00 per hr. $16.00 per hr.

I Love My Chocolate Company does not expect there to be any beginning or ending inventories of cocoa or sugar. At the end of the budget year, I Love My Chocolate Company had the following actual results:

Dark Chocolate Light Chocolate
Actual production (cases) 4,000 14,400
     Actual Price per Pound      Actual Pounds Purchased and Used
Cocoa $4.40 178,500
Sugar 0.55 235,600
Actual Labor Rate      Actual Labor Hours Used
Dark chocolate $15.60 per hr. 1,460
Light chocolate 16.40 per hr. 7,380

Required:

1. Prepare the following variance analyses for both chocolates and the total, based on the actual results and production levels at the end of the budget year:

     a. Direct materials price variance, direct materials quantity variance, and total variance.

     b. Direct labor rate variance, direct labor time variance, and total variance.

Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number. If there is no variance, enter a zero.

a. Direct materials price variance $
Direct materials quantity variance $
Total direct materials cost variance $
b. Direct labor rate variance $
Direct labor time variance $
Total direct labor cost variance $

2. The variance analyses should be based on the   amounts at   volumes. The budget must flex with the volume changes. If the   volume is different from the planned volume, as it was in this case, then the budget used for performance evaluation should reflect the change in direct materials and direct labor that will be required for the   production. In this way, spending from volume changes can be separated from efficiency and price variances.
Answer #1
Cocoa
calculation of direct material price variance:
= (Standard price per unit of material - Actual price per unit of material) × Actual quantity
= ($4.30/ib. - $4.40/ib ) × 178500 ib. = $17850 U F
Calculation of direct material quantity variance
=(standard quantity of material required for actual production - actual quantity used) × Standard price per unit
((12 ibs. X 4000 Unit)+(9ibsX14400)-178500)X $4.30= $3870 U F
For Sugar
calculation of direct material price variance:
= (Standard price per unit of material - Actual price per unit of material) × Actual quantity
= ($0.60 - $0.55 ) × 235600 = $11780 F
Calculation of direct material quantity variance
=(standard quantity of material required for actual production - actual quantity used) × Standard price per unit
((10 ibs. X 4000 Unit)+(14 ibsX14400)-235600)X $0.60= $3600 F
Total Material Price Variance= $17850 UF+$11780 F= $6070 UF
Total Material Quantity Variance= $3870 UF+ $3600 F=$270 UF
Total Material Cost Variance= $6070 UF+ $270 UF= $6340UF
Dark Choclate
Calculation of direct labor rate variance
= (Standard direct labor rate per hour - actual direct labor rate per hour) × Actual hours used
= ($16/hour - $15.60/hour) × 1460 Hours= $584 F
Calculation of direct labor efficiency variance:
= (standard hours required for actual production - actual hours used) × standard overhead recovery rate
= (0.40 Hour × 4200 Unit - 1460) × $16 = $3520 F
Light Choclate
Calculation of direct labor rate variance
= (Standard direct labor rate per hour - actual direct labor rate per hour) × Actual hours used
= ($16/hour - $16.40/hour) × 7380 Hours= $2952 UF
Calculation of direct labor efficiency variance:
= (standard hours required for actual production - actual hours used) × standard overhead recovery rate
= (0.50 Hour × 13800 Unit - 7380) × $16 = $7680 U F
Total LAbour Rate Variance= $584 F+ $2952 UF= $2368 UF
Total Labour Efficiency Variance= $3520F+$7680 UF=$4160 U F
Total Labour Cost Variance = $2368UF+ $4160 UF=$6528UF